In triangle STU, angle S is twice the measure of angle T and angle U is twice the measure of angle T increased by 20. Find the measure of all three angles
Answer by addingup(3677)

In every triangle, the sum of the angles equals 180
S+T+U = 180 (1)
and the problem says that:
S = 2T
U = 2T+20
Using these values for S and U, substitute in (1) above:
2T+T+2T+20 = 180
5T = 160
T = 32
Now let's find the others:
S = 2T = 2(32) = 64
U = 2T+20 = 2(32)+20 = 84
-----------------------------------
Check:
S+T+U = 180
64+32+84 = 180 Correct